40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11126     Entry time: Tue Mar 10 03:37:03 2015
Author: ericq 
Type: Update 
Category: LSC 
Subject: Locking efforts 

[Q, J]

Not much luck locking tonight; we made the RF transition to CARM numerous times, but it never lasted more than a minute or so. We were able to take a couple of loop and spectrum measurements as we transistioned. 

Here are some spectra showing the noise evolution of CARM_IN1 and DARM_IN1 as we start to transition CARM to RF. We did not manage to grab spectra while CARM was RF only; we can go back in the DQ to find some data. 

As we transition, our phase bubble is shrinking, which may explain our poor stability. On the following plot, I actually mistyped the legend. The cyan trace is ALL RF. I'm not sure why we have a 1/f^2 shape from 100->200Hz. 

[


We adjusted the pole compensation frequency by looking at REFL11/ALS during a CARM swept sine measurement, the -3db/-45degree point looked more like 80Hz. Strangely, the compensated REFL11 signal appears to lag the ALS signal around the UGF. Maybe this is a loop effect? 


In terms of practical improvements, I've written a script that reliably transitions from POX/POY IR lock to ALS CARM/DARM lock already on resonance. This is saving us a bunch of time. I've svn'd the new ALS script and the new carm_cm_up that uses it. 


We looked into the odd oplev behavior as well. We had earlier seen what looked like railed values on the FM output medm screen (which seemed unexpected for an AC coupled loop), but dataviewer showed it was actually ringing/railing at some 10+Hz as the oplev beam fell off the QPD. The ringing continues even after the quadrant values stop crossing zero, so I think it may be the filters themselves misbehaving. Why there is new behavior here is still beyond me. 


We lost a fair bit of time to a fussy mode cleaner tonight; there was a good 45 minute stretch where it refused to lock for more than a minute or so, the PC drive angriliy never falling below 5. The thing I changed when it started working was using the fast C1:IOO-MC_F channel instead of the slow C1:IOO-MC_FAST_MON as a readback for the FSS input offset; oddly there is a DC difference between the two. This has resulted in a FSS offset of ~4.2, whereas it was previously ~1.8. After this change, the PC drive fell to ~1.0 levels, and the IMC has been mostly ok. 


Given our problems stabilizing the RF lock, we attempted to give the FOOL path a shot, since we now had a better idea of the neccesary REFL11 gain. In short, no luck. Every attempt to use some RF signal just disturbed the lock further. We didn't really pursue it too much after a couple of attempts showed little promise. 

Attachment 1: 2015-03-10_rfCarmOLG.png  208 kB  Uploaded Tue Mar 10 04:39:56 2015  | Hide | Hide all
2015-03-10_rfCarmOLG.png
Attachment 2: 2015-03-10_rfTransitions.png  404 kB  Uploaded Tue Mar 10 04:42:58 2015  | Hide | Hide all
2015-03-10_rfTransitions.png
ELOG V3.1.3-